OpenStudy (anonymous):

Suppose a runner takes 45 min to run a route at 8 mi/h at the beginning of training season. By the end of training season, she can run the same route in 38 min. What is her speed at the end of training season?

OpenStudy (mathmale):

We're looking for a measure of speed. Note that speed = distance/time. The distance remains the same, but the length of time required by the runner to run that distance decreases. What is the distance run? What is the time required for the runner to run this distance by the end of the training season? Now, using that formula for speed, calculate the speed.

OpenStudy (mathmale):

@student001 : Thanks for your input. Would you please help the other person along the road to finding his or her own answer, and NOT provide the answer yourself.

OpenStudy (anonymous):

Mathmale just because you're the monirater or whatever. doesnt give you the right to dlete ppl's answers

OpenStudy (anonymous):

*Moderator

OpenStudy (mathmale):

Note: You'll have to use the info given: 45 minutes, 8 miles/hour to find out how far the runner has to run. Best to start with that.

OpenStudy (mathmale):

Hint: how many minutes in an hour? 45 minutes is what fraction of an hour?

OpenStudy (anonymous):

Like 3/4 ?

OpenStudy (mathmale):

Very good. So, if this person runs at 8 mph for 3/4 hour, how far does he/she run?

OpenStudy (anonymous):

Idk

OpenStudy (mathmale):

distance = rate * time. So, distance = (8 miles/hour)*(3/4 hour) = ? Come on...give this a try.

OpenStudy (mathmale):

\[Hint:\frac{ 8.miles }{ hour }*\frac{ 3.hours }{ 4 }=?\]

OpenStudy (anonymous):

Sorry if my responses are late, my internet is really slow today.

OpenStudy (mathmale):

\[\frac{ 8 }{ 1 }*\frac{ 3 }{ 4 }=?\]

OpenStudy (anonymous):

Would it be like this \[\frac{ 8 }{ 1 }\times \frac{ 3 }{ 4 }=\frac{ 24 }{ 4 }?\]

OpenStudy (mathmale):

Sure! But it'd be easier if you were to divide that 8 by that 4 first, obtaining 2*3=6. So: The runner runs 6 miles. Supposing it takes her 38 minutes to run that far ... or 38/60 hours, please calculate her end-of-season speed.